# Sample Compression for Continual Learning

### v1

@@ Abstract
Continual learning algorithms aim to learn from a sequence of tasks, making the training distribution non-stationary. The majority of existing continual learning approaches in the literature rely on heuristics and do not provide learning guarantees for the continual learning setup. In this paper, we present a new method called ‘Continual Pick-to-Learn’ (CoP2L), which is able to retain the most representative samples for each task in an efficient way. The algorithm is adapted from the Pick-to-Learn algorithm, rooted in the sample compression theory. This allows us to provide high-confidence upper bounds on the generalization loss of the learned predictors, numerically computable after every update of the learned model. We also empirically show on several standard continual learning benchmarks that our algorithm is able to outperform standard experience replay, significantly mitigating catastrophic forgetting.
Machine Learning, ICML
@@ 1 Introduction
A common assumption in traditional machine learning is that the underlying data distribution does not evolve with time. In Continual Learning (De Lange et al., 2021; Wang et al., 2024), the goal is to develop machine learning algorithms that are able to learn under settings where this assumption is replaced by a set-up where the model is trained on an evolving training data distribution, in such a way that samples are revealed one task at a time.
However, when neural networks are trained on evolving data distributions, networks tend to forget the training from earlier tasks. This fact is known as the catastrophic forgetting (McCloskey & Cohen, 1989; French, 1999; Goodfellow et al., 2014). In order to cope with forgetting, various methodologies have been developed, such as the regularization based approaches (e.g., Kirkpatrick et al., 2017), architectural approaches (e.g., Rusu et al., 2016), or rehearsal based approaches (e.g., Rolnick et al., 2019).
It has been shown on various applications that rehearsal-based approaches typically perform very competitively compared to the other methods in terms of performance and forgetting (Libera et al., 2023; Mai et al., 2020; Chaudhry et al., 2019; Aljundi et al., 2019). However, the selection of the replay buffer typically remains a heuristic, and precise upper bounds for the generalization error are not available.
@@
Our proposed continual learning scheme builds upon the Pick-to-Learn algorithm (P2L) of Paccagnan et al. (2024). P2L is a meta algorithm that selects a compression set, which amounts to express the dataset in terms of ‘core’ examples. This meta-algorithm was developed specifically with sample compression theory (Floyd & Warmuth, 1995; Campi & Garatti, 2023; Laviolette et al., 2005) in mind. By finding a small subset of the data such that a predictor learned on this data achieves low error on the whole training set, Pick-to-Learn enables us to compute tight upper bounds for the generalization error of the learned predictor.
@@
In this paper, we propose ‘Continual Pick-to-Learn‘ (CoP2L), an algorithm that leverages the sample compression theory to intelligently select the training data from earlier tasks to mitigate forgetting. We derive high-confidence upper bounds on the generalization error for each task simultaneously, estimated directly from the training set. We also empirically show that CoP2L mitigates the forgetting significantly, and outperforms the standard replay baseline significantly in the small replay buffer size settings. Our contribution overall can be summarized as follows :
@@
  • We propose the algorithm Continual Pick-to-Learn (CoP2L), that integrates the sample compression theory within the continual learning setup. To the best of our knowledge, we are the first to integrate the theoretical results from sample compression to continual learning.
@@
  • CoP2L is able to provide high confidence non-trivial upper bounds for the generalization error by means of the sample compression theory. We experimentally showcase that the bounds are numerically computable, have non-trivial values and follow the general error trends observed on the test set. Therefore, they can actually be used as risk certificates on the model behavior on the learned tasks.
@@
  • We experimentally show that CoP2L significantly mitigates forgetting, while also being able to obtain better performance compared to vanilla replay, in the settings where the replay buffer is small.
@@ 1.1 Related Work
Practical Continual learning
In the continual learning literature, various practical approaches have been developed to mitigate forgetting. Broadly, the approaches could be divided into three categories which include regularization approaches (Kirkpatrick et al., 2017; Zenke et al., 2017; Li & Hoiem, 2018; Aljundi et al., 2018), architecture-based approaches (Rusu et al., 2016; Mallya et al., 2018; Aljundi et al., 2017; Mallya & Lazebnik, 2018; Wang et al., 2022) and rehearsal based approaches (Rolnick et al., 2019; Chaudhry et al., 2019; Shin et al., 2017; Rebuffi et al., 2017; Buzzega et al., 2020). In this paper, we use experience replay as a baseline as it is shown in the literature to be a very strong continual learning baseline to mitigate forgetting (Libera et al., 2023).
Theoretical Continual Learning
There have been several works in theoretical continual learning in the recent years. For example, Bennani & Sugiyama (2020); Doan et al. (2021) study the generalization error obtained with Neural Tangent Kernel (NTK) models when the model is trained with Orthogonal Gradient Descent (Farajtabar et al., 2020).
Yin et al. (2020) provides generalization bounds for regularization-based continual learning methods. Asanuma et al. (2021) studies the impact of task similarity in the student-teacher setup. Li et al. (2022) provides a sample complexity bound in the context of continual representation learning for an adapted PackNet (Mallya & Lazebnik, 2018). In Chen et al. (2022), the authors provide a theoretical bound on memory using PAC learning theory. These works however do not provide bounds that are computable in practice.
A line of work also investigates the generalization of continual learning models when linear models are employed. In Goldfarb & Hand (2023), the authors investigate the effect of overparametrized linear models. In Evron et al. (2022); Lin et al. (2023), the authors analyze the generalization behavior of linear regression models, even though Lin et al. (2023) conjectures that their results could be extended to neural networks as well.
Different from all the listed works above, in this paper, we provide a practical non-trivial generalization bound that is computed on the training set, and predicts the test error for each task separately. Furthermore, also unlike the approaches above, our bound is applicable to any neural network architecture and does not have limitations on the continual learning setup that is employed.
Sample Compression
The sample compress theory has been shown effective for providing tight generalization bounds (Laviolette et al., 2005; Marchand et al., 2003; Marchand & Shawe-Taylor, 2002; Marchand & Sokolova, 2005; Shah, 2007). However, most sample compress approaches are limited to low-complexity models. A notable exception is the method Pick-to-Learn (P2L), which successfully provides guarantees for deep neural networks (Campi & Garatti, 2023; Paccagnan et al., 2024). In this paper, we adapt the key ideas introduced in the sample compression theory to the continual learning case, which enables us to provide an upper bound on the test error from training samples.
@@ 2 Background
2.1 Continual Learning
Notation.
Let us consider a series of task distributions ${\mathcal{D}}_{1},{\mathcal{D}}_{2},\ldots,{\mathcal{D}}_{T}$ on an input-output space ${\mathcal{X}}\times{\mathcal{Y}}$, where the amount of tasks $T\in{\mathbb{N}}^{*}$ can be (countably) infinite. We are given a predictive model $f_{\theta}:{\mathcal{X}}\to{\mathcal{Y}}$ (e.g., a neural network architecture), with learnable parameters $\theta\in\Theta$. From randomly initialized parameters $\theta_{0}$, the aim of the continual learning process at step $t\in\{1,\ldots,T\}$ is to update the parameters $\theta_{t-1}$ into $\theta_{t}$ by learning from a new task sample $S_{t}\sim{\mathcal{D}}_{t}$.
We denote the training set at task $t$ as $S_{t}=\{({\mathbf{x}}_{t,i},y_{t,i})\}_{i=1}^{n_{t}}$, with ${\mathbf{x}}_{t,i}\in{\mathcal{X}}$ and $y_{t,i}\in{\mathcal{Y}}_{t}\subseteq{\mathcal{Y}}$; that is, the same input space ${\mathcal{X}}$ is shared among all tasks, and the output space ${\mathcal{Y}}$ may be the union of several task specific output spaces ${\mathcal{Y}}_{t}$.
Given a loss function $\ell:\Theta\times{\mathcal{X}}\times{\mathcal{Y}}\to{\mathbb{R}}^{+}$, we want the last updated predictor $f_{\theta_{T}}$ to maintain a low generalization loss on all observed tasks $t\in\{1,\ldots,T\}$:
@@ 2.1 Continual Learning
The challenge of continual learning lies in the fact that the learner observes the task datasets $S_{1},S_{2},\ldots,S_{t},\ldots,S_{T}$ sequentially, and we assume that the system cannot keep all observed data in memory. Nevertheless, at task $t$, we would like the system to perform well on all tasks, including the previous ones $11,\dots,t-$ 1. However, simply updating the parameters $\theta_{t-1}$ learned from a previous tasks to optimize the loss ^$\widehat{\mathcal{L}}_{S_{t}}(\theta_{t})$ on a current task $t$ would lead to catastrophic forgetting (French, 1999). Therefore, as indicated in Section 1.1, numerous empirical methods have been developed to mitigate forgetting. A simple yet very effective strategy is to simply store a small subset of data from the earlier tasks, coined as the replay buffer, and to include them in the objective function of the experience replay strategy for learning task $t$:
@@
where the second term is the loss on the replay buffer ${\mathcal{B}}$. In our experiments, we use this method as the baseline as it’s shown to be a simple and very effective continual learning strategy (Libera et al., 2023; Mai et al., 2020).
2.2 Sample Compression Theory
Given a dataset $S=\{(\boldsymbol{x}_{i},y_{i})\}_{i=1}^{n}$, a family of learnable parameters $\Theta$ and a learning algorithm $A$ such that $A(S)\in\Theta$, sample compression theory provides generalization bounds for any predictor $A(S)$ on the condition that the predictor can be provably represented as a function of a small subset of the dataset, called the compression set, and an additional source of information, called the message. If this is the case, we call $A(S)$ a sample-compressed predictor.
The compression set is denoted $S^{\operatorname{\mathbf{i}}}$ and is parameterized by a vector of indices $\operatorname{\mathbf{i}}\subset\{1,\ldots,n\}$. Each vector $\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)$, with $\operatorname{\mathscr{P}}(n)$ the set of all the possible subsets of $\{1,\ldots,n\}$, is ordered such that we have
@@ 2.2 Sample Compression Theory
The compression set is defined such that
$S^{\operatorname{\mathbf{i}}}=S^{(i_{1},\ldots,i_{\operatorname{|\mathbf{i}|}})}\!=\{(\boldsymbol{x}_{i_{1}},y_{i_{1}}),\ldots,(\boldsymbol{x}_{i_{\operatorname{|\mathbf{i}|}}},y_{i_{\operatorname{|\mathbf{i}|}}})\}\!\subseteq S.$ (5)
We define the complement vector $\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}$ such that $\operatorname{\mathbf{i}}\cap\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}=\emptyset$ and $\operatorname{\mathbf{i}}\cup\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}=\{1,\ldots,n\}$. Thus, we have the complement set $S_{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}=S\setminus S_{\operatorname{\mathbf{i}}}$ and $|\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}|=n-\operatorname{|\mathbf{i}|}$.
@@
Given a learning algorithm $A$ and a dataset $S$, to prove that the predictor $A(S)$ is a sample-compressed one, we need to define two functions: a compression function and a reconstruction function. The compression function $\mathscr{C}:\bigcup_{1\leq n\leq\infty}(\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}})^{n}\to\bigcup_{m\leq n}(\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}})^{m}\times\bigcup_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\mathscr{M}(\operatorname{\mathbf{i}})$ outputs a compression set and a message that are used to represent $A(S)$. Given a dataset $S$, we have $\mathscr{C}(S)=(S^{\operatorname{\mathbf{i}}},\mu)$. Then, the reconstruction function $\operatorname{\mathscr{R}}:\bigcup_{m\leq n}(\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}})^{m}\times\bigcup_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\mathscr{M}(\operatorname{\mathbf{i}})\to\Theta$ is defined such that $A(S)=\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},\mu)$. Both the reconstruction function and the compression function must be deterministic and data-independent.
The forthcoming results rely on a probability distribution over the set of sample-compressed predictors $\overline{\Theta}\subseteq\Theta$. For any sample-compressed predictor $\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},\mu)$, this data-independent distribution is expressed as $P_{\overline{\Theta}}\quantity(\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},\mu))=P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}})P_{\mathscr{M}(\operatorname{\mathbf{i}})}(\mu)$, with $P_{\operatorname{\mathscr{P}}(n)}$ a probability distribution over $\operatorname{\mathscr{P}}(n)$ and $P_{\mathscr{M}(\operatorname{\mathbf{i}})}$ a probability distribution over $\mathscr{M}(\operatorname{\mathbf{i}})$. Following the work of Marchand & Sokolova (2005), we consider $P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}})=\smallmatrixquantity(n\\ \operatorname{|\mathbf{i}|})^{-1}\zeta(\operatorname{|\mathbf{i}|})$, with $\zeta(k){=}\tfrac{6}{\pi^{2}}(k+1)^{-2}$. Given a vector of indices $\operatorname{\mathbf{i}}$, we consider $P_{\mathscr{M}(\operatorname{\mathbf{i}})}$ to be a uniform distribution over the messages.
We now present the sample compression bound of Bazinet et al. (2025). We use the shorthand notation ^$\operatorname{\widehat{\mathcal{L}}}_{S}^{\,{}_{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}}(\cdot)$ to denote the empirical loss on the training sample that don’t belong to the compression set: $S^{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}=S\setminus S^{\mathbf{i}}$; this is mandatory to obtain an unbiased estimate of the loss, as the reconstructed predictor $\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},\mu)$ relies on $S^{\mathbf{i}}$.
Theorem 2.1.
@@
$\mathrm{kl}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,{}_{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}}\quantity(\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},\mu)),\mathcal{L}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},\mu)))\leq\epsilon(n,\operatorname{\mathbf{i}},\mu,\delta)$
with $\mathrm{kl}(q,p)=q\ln\frac{q}{p}+(1-q)\ln\frac{1-q}{1-p}$ the binary Kullback-Leibler divergence and
$\epsilon(n,\operatorname{\mathbf{i}},\mu,\delta)=\frac{1}{n-\operatorname{|\mathbf{i}|}}\quantity[\log\!\matrixquantity(n\\ \operatorname{|\mathbf{i}|})\!\!+\!\log\!\quantity(\frac{2\sqrt{n-\operatorname{|\mathbf{i}|}}}{\zeta(\operatorname{|\mathbf{i}|})P_{\mathscr{M}(\operatorname{\mathbf{i}})}(\mu)\delta})].$
Given the parameters $\theta$ of a reconstructed predictor, i.e., $\theta{=}\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},\mu)$, Theorem 2.1 states that the $\mathrm{kl}$-discrepancy between the true risk ^$\operatorname{\widehat{\mathcal{L}}}_{S}^{\,{}_{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}}(\theta)$ and the empirical risk $\mathcal{L}_{\operatorname{\mathcal{D}}}(\theta)$ is upper bounded by the complexity term $\epsilon(n,\operatorname{\mathbf{i}},\mu,\delta)$. From this result, one can compute an upper bound on the true risk $\mathcal{L}_{\operatorname{\mathcal{D}}}(\theta)$ by inverting the Kullback-Leibler divergence. Indeed, Theorem 2.1 can be rewritten as
$\mathcal{L}_{\operatorname{\mathcal{D}}}\quantity(\theta)\leq\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,{}_{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}}\quantity(\theta),\epsilon(n,\operatorname{\mathbf{i}},\mu,\delta))$ (6)
with
$\mathrm{kl}^{-1}(q,\varepsilon)\ =\ \operatorname*{arg\,sup}_{0\leq p\leq 1}\{\mathrm{kl}(q,p)\leq\varepsilon\}.$ (7)

### v3

# Sample Compression for Self-Certified Continual Learning

@@ Abstract
Continual learning algorithms aim to learn from a sequence of tasks, making the training distribution non-stationary. The majority of existing continual learning approaches in the literature rely on heuristics and do not provide learning guarantees. In this paper, we present a new method called Continual Pick-to-Learn (CoP2L), which is able to retain the most representative samples for each task in an efficient way. CoP2L combines the Pick-to-Learn algorithm (rooted in the sample compression theory) and the experience replay continual learning scheme. This allows us to provide non-vacuous upper bounds on the generalization loss of the learned predictors, numerically computable after each task. We empirically evaluate our approach on several standard continual learning benchmarks across Class-Incremental, Task-Incremental, and Domain-Incremental settings. Our results show that CoP2L is highly competitive across all setups, often outperforming existing baselines, and significantly mitigating catastrophic forgetting compared to vanilla experience replay in the Class-Incremental setting. It is possible to leverage the bounds provided by CoP2L in practical scenarios to certify the predictor reliability on previously learned tasks, in order to improve the trustworthiness of the continual learning algorithm.
@@ 1 Introduction
A common assumption in traditional machine learning is that the underlying data distribution does not evolve with time. In Continual Learning de2021continual ; wang2023comprehensive , the goal is to develop machine learning algorithms that are able to learn under settings where this assumption is replaced by a set-up where the model is trained on an evolving training data distribution, in such a way that samples are revealed one task at a time. However, when neural networks are trained on evolving data distributions, networks tend to forget the training from earlier tasks. This fact is known as the catastrophic forgetting McCloskey1989CatastrophicII ; forgetting_french ; goodfellow2013empirical . In order to cope with forgetting, various methodologies have been developed, such as regularization based approaches ewc_kirkpatrick ; zenke2017continual ; cl_li ; mas , architectural approaches pnn_rusu ; mallya2018piggyback ; expert_aljundi ; packnet_mallya ; l2p_wang , or rehearsal based approaches rolnick2019experience ; chaudhry2019efficient ; replay_shin ; icarl_rebuffi ; der . It has been shown on various applications that rehearsal-based approaches typically perform very competitively compared to the other methods in terms of performance and forgetting DellaLibera2023CLMASRAC ; Mai2020BatchlevelER ; chaudhry2019efficient ; aljundi2019onlinecontinuallearningmaximally . However, the selection of the replay buffer typically remains a heuristic, and precise upper bounds for the generalization error are not available.
@@
Our proposed continual learning scheme builds upon the Pick-to-Learn algorithm (P2L) of Paccagnan et al. paccagnan_pick_learn_2023 . P2L is a meta-algorithm that selects a compression set, which amounts to expressing the dataset in terms of ‘core’ examples. This meta-algorithm was developed specifically with sample compression theory in mind floyd1995sample ; campi2023compression . By finding a small subset of the data such that a predictor learned on this data achieves low error on the whole training set, Pick-to-Learn enables us to compute tight upper bounds for the generalization error of the learned predictor. This ability of a learning algorithm to simultaneously output a predictor and its risk certificate has been coined as self-certified learning in the literature freund98selfbounding ; perezortiz2021tighter ; marks2025pick2learn-gp . This certification is arguably an asset to increase the trustworthiness towards machine learning systems.
@@
In this paper, we propose Continual Pick-to-Learn (CoP2L), an algorithm that leverages sample compression theory to intelligently select the training data from earlier tasks to mitigate forgetting. We derive high-confidence upper bounds on the generalization error for each task simultaneously, estimated directly from the training set. We also empirically show that CoP2L mitigates the forgetting significantly and outperforms the standard replay baseline significantly in the small replay buffer size settings. Our contributions can be summarized as follows:
@@
(1) We propose the algorithm Continual Pick-to-Learn (CoP2L), that integrates the sample compression theory within the continual learning setup. To the best of our knowledge, we are the first to integrate the theoretical results from sample compression to continual learning.
@@
(2) CoP2L is able to provide high confidence non-trivial upper bounds for the generalization error by means of the sample compression theory. We experimentally showcase that the bounds are numerically computable, have non-trivial values and follow the general error trends observed on the test set. Therefore, they can actually be used as risk certificates on the model behavior on the learned tasks, improving the trustworthiness of the continually learnt model.
@@
(3) We experimentally show with a wide range of experiments that CoP2L significantly mitigates forgetting, while also being able to obtain better or comparable performance compared with several strong continual learning baselines in most cases.
@@ 1.1 Related Work
Continual learning approaches.  In the continual learning literature, various practical approaches have been developed to mitigate forgetting. Broadly, the approaches could be divided into three categories, which include regularization approaches ewc_kirkpatrick ; zenke2017continual ; cl_li ; mas , architecture-based approaches pnn_rusu ; mallya2018piggyback ; expert_aljundi ; packnet_mallya ; l2p_wang and rehearsal based approaches rolnick2019experience ; chaudhry2019efficient ; replay_shin ; icarl_rebuffi ; der .
Continual learning generalization bounds.  In recent years, generalization bounds were proposed to study the asymptotic behavior of various continual learning schemes, for example using linear models Evron2022HowCC ; Goldfarb2022AnalysisOC ; pmlr-v202-lin23f , Neural Tangent Kernel (NTK) models trained with Orthogonal Gradient Descent AbbanaBennani2020GeneralisationGF ; Doan2020ATA ; farajtabar2019orthogonalgradientdescentcontinual , regularization-based methods yin2021optimizationgeneralizationregularizationbasedcontinual , or PackNet packnet_mallya inspired continual representation learning Li2022ProvableAE . Learning algorithms with convergence guarantees are were also proposed under data assumptions, such that all target tasks lie in a same space obtainable from a (linear or non-linear) combination of some input features cao2022lifelong .
Different from all the listed works above, in this paper, we provide a generalization bound that is computed from the training set. Risk certificates on the test error are provided simultaneously for each task. Furthermore, also unlike the approaches above, our non-asymptotic bound is applicable to any neural network architecture, under minimal data assumptions, and does not have limitations on the continual learning setup that is employed.
Sample compression.  The sample compression theory has been shown effective for providing tight generalization bounds shah_margin-sparsity_2005 ; marchand2003set ; marchand2002set ; marchand2005learning ; shah_sample_2007 . However, most sample compress approaches are limited to low-complexity models. A notable exception is the method Pick-to-Learn (P2L), which successfully provides guarantees for deep neural networks campi2023compression ; bazinet2024 and Gaussian Processes marks2025pick2learn-gp . In this paper, we adapt the key ideas introduced in the sample compression theory to the continual learning case, which enables us to provide an upper bound on the test error from training samples.
@@ 2 Background
Continual learning.  Let us consider a series of task distributions ${\mathcal{D}}_{1},{\mathcal{D}}_{2},\ldots,{\mathcal{D}}_{T}$ on an input-output space ${\mathcal{X}}\times{\mathcal{Y}}$, where the amount of tasks $T\in{\mathbb{N}}^{*}$ can be (countably) infinite. We are given a predictive model $f_{\theta}:{\mathcal{X}}\to{\mathcal{Y}}$ (e.g., a neural network architecture), with learnable parameters $\theta\in\Theta$. From randomly initialized parameters $\theta_{0}$, the aim of the continual learning process at step $t\in\{1,\ldots,T\}$ is to update the parameters $\theta_{t-1}$ into $\theta_{t}$ by learning from a new task sample $S_{t}\sim{\mathcal{D}}_{t}$. We denote the training set at task $t$ as $S_{t}=\{({\mathbf{x}}_{t,i},y_{t,i})\}_{i=1}^{n_{t}}$, with ${\mathbf{x}}_{t,i}\in{\mathcal{X}}$ and $y_{t,i}\in{\mathcal{Y}}_{t}\subseteq{\mathcal{Y}}$; that is, the same input space ${\mathcal{X}}$ is shared among all tasks, and the output space ${\mathcal{Y}}$ may be the union of several task specific output spaces ${\mathcal{Y}}_{t}$. Given a loss function $\ell:\Theta\times{\mathcal{X}}\times{\mathcal{Y}}\to{\mathbb{R}}^{+}$, we want the last updated predictor $f_{\theta_{T}}$ to maintain a low generalization loss on all observed tasks $t\in\{1,\ldots,T\}$:
@@
The challenge of continual learning lies in the fact that the learner observes the task datasets $S_{1},S_{2},\ldots,S_{t},\ldots,S_{T}$ sequentially, and we assume that the system cannot keep all observed data in memory. Nevertheless, at task $t$, we would like the system to perform well on all tasks, including the previous ones $11,\dots,t-$ 1. However, simply updating the parameters $\theta_{t-1}$ learned from a previous tasks to optimize the loss ^$\widehat{\mathcal{L}}_{S_{t}}(\theta_{t})$ on a current task $t$ would lead to catastrophic forgetting forgetting_french . Therefore, as indicated in Section 1.1, numerous empirical methods have been developed to mitigate forgetting. A simple yet very effective strategy is to simply store a small subset of data from the earlier tasks, coined as the replay buffer, and to include them in the objective function of the experience replay strategy for learning task $t$:
@@
where the second term is the loss on the replay buffer ${\mathcal{B}}$. In our experiments, we use this method as a baseline as it’s shown to be a simple and very effective continual learning strategy DellaLibera2023CLMASRAC ; Mai2020BatchlevelER .
Sample compression theory.  Given a training dataset $S=\{(\boldsymbol{x}_{i},y_{i})\}_{i=1}^{n}\in(\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}})^{n}$, a family of learnable parameters $\Theta$ and a learning algorithm $A$ such that $A(S)\in\Theta$, sample compression theory provides generalization bounds for any predictor $f_{\theta}$, with $\theta=A(S)$, on the condition that $f_{\theta}$ can be provably represented as a function of a small subset of the training dataset $S$, called the compression set, and an additional source of information, called the message. If this is the case, we call $f_{\theta}$ a sample-compressed predictor. The compression set is denoted $S^{\operatorname{\mathbf{i}}}$ and is parameterized by a vector of indices $\operatorname{\mathbf{i}}\subset\{1,\ldots,n\}$. Each vector $\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)$, with $\operatorname{\mathscr{P}}(n)$ the set of all the possible subsets of $\{1,\ldots,n\}$, is ordered such that
@@
The compression set is defined such that $S^{\operatorname{\mathbf{i}}}=S^{(i_{1},\ldots,i_{\operatorname{|\mathbf{i}|}})}\!=\{(\boldsymbol{x}_{i_{1}},y_{i_{1}}),\ldots,(\boldsymbol{x}_{i_{\operatorname{|\mathbf{i}|}}},y_{i_{\operatorname{|\mathbf{i}|}}})\}\!\subseteq S.$ We define the complement vector $\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}$ such that $\operatorname{\mathbf{i}}\cap\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}=\emptyset$ and $\operatorname{\mathbf{i}}\cup\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}=\{1,\ldots,n\}$. Thus, we have the complement set $S^{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}=S\setminus S^{\operatorname{\mathbf{i}}}$ and $|\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}|=n-\operatorname{|\mathbf{i}|}$.
@@
Given learned parameters $\theta=A(S)$, to prove that the predictor $f_{\theta}$ is a sample-compressed one, we need to define two functions: a compression function and a reconstruction function. The compression function $\mathscr{C}:\boldsymbol{\cup}_{1\leq n\leq\infty}(\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}})^{n}\to\boldsymbol{\cup}_{m\leq n}(\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}})^{m}\times\boldsymbol{\cup}_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\mathscr{M}(\operatorname{\mathbf{i}})$ outputs a compression set and a message. Then, the reconstruction function $\operatorname{\mathscr{R}}:\boldsymbol{\cup}_{m\leq n}(\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}})^{m}\times\boldsymbol{\cup}_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\mathscr{M}(\operatorname{\mathbf{i}})\to\Theta$ is defined such that $A(S)=\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},\mu)$. The forthcoming results rely on a probability distribution over the set of sample-compressed predictors $\overline{\Theta}\subseteq\Theta$. For any sample-compressed predictor $\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},\mu)$, this data-independent distribution is expressed as $P_{\overline{\Theta}}\quantity(\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},\mu))=P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}})P_{\mathscr{M}(\operatorname{\mathbf{i}})}(\mu)$, with $P_{\operatorname{\mathscr{P}}(n)}$ a probability distribution over $\operatorname{\mathscr{P}}(n)$ and $P_{\mathscr{M}(\operatorname{\mathbf{i}})}$ a probability distribution over $\mathscr{M}(\operatorname{\mathbf{i}})$. Following common practices marchand2005learning ; bazinet2024 , we consider hereunder $P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}})={n\choose\operatorname{|\mathbf{i}|}}^{{}_{-1}}\zeta(\operatorname{|\mathbf{i}|})$, with $\zeta(k){=}\tfrac{6}{\pi^{{}_{2}}}(k+1)^{-2}$. Also, given a vector of indices $\operatorname{\mathbf{i}}$, we consider $P_{\mathscr{M}(\operatorname{\mathbf{i}})}$ to be a uniform distribution over the messages.
We now present a recent sample compression generalization guarantee bazinet2024 on which we build our theoretical framework. We use the shorthand notation ^$\operatorname{\widehat{\mathcal{L}}}_{S}^{\,{}_{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}}(\cdot)$ to denote the empirical loss on the training sample that do not belong to the compression set: $S^{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}=S\setminus S^{\mathbf{i}}$; this is mandatory to obtain an unbiased estimate of the loss, as the reconstructed predictor $\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},\mu)$ relies on $S^{\mathbf{i}}$.
Theorem 2.1 (bazinet2024 ).
@@
$\mathcal{L}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},\mu))\leq\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,{}_{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}}\quantity(\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},\mu)),\frac{1}{n-\operatorname{|\mathbf{i}|}}\quantity[\log\!\matrixquantity(n\\ \operatorname{|\mathbf{i}|})\!\!+\!\log\!\quantity(\frac{2\sqrt{n-\operatorname{|\mathbf{i}|}}}{\zeta(\operatorname{|\mathbf{i}|})P_{\mathscr{M}(\operatorname{\mathbf{i}})}(\mu)\delta})])\,,$
with $\mathrm{kl}(q,p)=q\ln\frac{q}{p}+(1-q)\ln\frac{1-q}{1-p}$ the binary Kullback-Leibler divergence and its inverse
$\mathrm{kl}^{-1}(q,\varepsilon)=\mbox{\rm argsup}_{p\in[0,1]}\left\{\mathrm{kl}(q,p)\leq\varepsilon\right\}.$ (5)